OpenStudy (anonymous):

Which explains why the sequence 81, 3, , 1/9… is arithmetic or geometric? • The sequence is arithmetic because it decreases by a factor of . • The sequence is geometric because it decreases by a factor of . • The sequence is arithmetic because it decreases by a factor of 3. • The sequence is geometric because it decreases by a factor of 3. My answer is C check my work

OpenStudy (welshfella):

divide the second term by the first and the third by the second

OpenStudy (welshfella):

3/81 = ? 3 / 1/9 = ?

OpenStudy (welshfella):

sorry thats 1/9 / 3

OpenStudy (welshfella):

its geometric because the next term is found by 3/81 = 1/27

OpenStudy (anonymous):

so D